OpenStudy (anonymous):

Help with his example? Find f(x) and g(x) such that h(x)=(f∘g)(x) h(x)=9/x-6 Suppose that g(x)=x-6 F(x)=

OpenStudy (anonymous):

\[ h(x)=9/x-6 = 9/g(x) \]

OpenStudy (anonymous):

\[ h(x) = f(g(x)) = 9/g(x) \implies f(x) = 9/x \]

OpenStudy (anonymous):

This is all about patter recognition.

OpenStudy (anonymous):

@wio Thank you! Your breakdown was easier to understand than my book. I think I can tackle some of these now.